### v1

# Generation of mechanical cat-like states via optomagnomechanics

Abstract
We propose an optomagnomechanical approach for preparing a cat-like superposition state of mechanical motion. Our protocol consists of two steps and is based on the magnomechanical system where the magnetostrictively induced displacement further couples to an optical cavity mode via radiation pressure. We first prepare a squeezed mechanical state by driving the magnomechanical system with a two-tone microwave field. We then switch off the microwave drives and send a weak red-detuned optical pulse to the optical cavity to weakly activate the optomechanical anti-Stokes scattering. We show that $k$ phonons can be subtracted from the prepared squeezed state, conditioned on the detection of $k$ anti-Stokes photons from the cavity output field, which prepares the mechanical motion in a cat-like state. The work provides a new avenue for preparing mechanical superposition states by combining opto- and magnomechanics and may find applications in the study of macroscopic quantum states and the test of collapse theories.
I Introduction
One distinctive feature of quantum mechanics is embodied by the fact that it allows the linear superposition of quantum states. As one special type of superposition states, the cat state refers to a quantum state composed of a superposition of two coherent states with opposite phases. Cat states were routinely generated in a variety of microscopic systems, including trapped ions CM96 ; DL05 , microwave photons BV13 optical photons KH15 ; AO06 ; JSN06 , atomic ensembles AO19 , and hybrid atom-light systems BH19 . However, preparing such states in macroscopic systems, e.g., a massive mechanical oscillator, is extremely challenging. Impressively, a recent experiment has demonstrated the so-far largest cat state of a bulk acoustic-wave resonator via the manipulation of a superconducting qubit MB23 . Apart from this system, proposals IS20 ; Tan20 indicate that the optomechancial system MA14 is also a promising system to prepare cat-like states of a macroscopic mechanical oscillator.
In recent years, the magnomechanical system Zuo has attracted much attention due to its potential for preparing macroscopic quantum states JL18 ; JL19b , quantum sensing Jing24 , and its rich nonlinearities, such as magnon-phonon cross-Kerr effect RCS22 and magnonic frequency combs Xiong23 ; Dong23 . For a large-size magnetic sample, e.g., an yttrium-iron-garnet (YIG) sphere, the magnetostrictive interaction is a dispersive type, which couples magnon excitations to the deformation displacement. Further coupling the magnomechanical displacement to an optical cavity via radiation pressure forms a hybrid optomagnomechanical (OMM) system Fan ; Dong22 ; QST23 , which has been proven to be a promising system to realize the optical readout of magnonic states QST23 and generate microwave-optics LPR23 ; LPR25 ; Xiong25 ; Zhang25 and magnon-atom entanglement FanPRA23 ; Di24 .
Here, we present a scheme for preparing a mechanical cat-like state based on the OMM system. We take the approach of subtracting excitations from the squeezed vacuum state MD97 ; AB07 . Specifically, our scheme consists of two steps: $i$) Preparing a mechanical squeezed state in the magnomechanical subsystem by applying a two-tone driving field to the magnon mode; $ii$) Subtracting phonons from the generated squeezed state by sending a weak red-detuned optical pulse to the optomechanical cavity. The detection of $k$ anti-Stokes photons in the cavity output field heralds a $k$-phonon-subtracted squeezed state, corresponding to a cat-like state of mechanical motion.
The paper is organized as follows. In Sec. II, we introduce the OMM system and show how to generate a mechanical squeezed state in the magnomechanical subsystem. In Sec. III, we show how to subtract $k$ phonons from the mechanical squeezed state in the optomechanical subsystem via the detection of the scattered anti-Stokes photons, and present the results of the mechanical cat-like states. Finally, we summarize our findings in Sec. IV.
II Magnomechanical SQUEEZING OF MECHANICAL motion
Figure 1: (a) Two-step protocol adopted to generate cat-like states of mechanical motion in an OMM system. First step: the magnon mode is driven by two microwave fields at the frequencies $\omega_{\pm}=\omega_{m}\pm\omega_{b}$ to prepare a mechanical squeezed state. Second step: a weak red-detuned optical pulse is sent to the optomechanical cavity to subtract phonons, conditioned on the detection of anti-Stokes photons in the cavity output field. (b) Frequencies of the magnomechanical subsystem and two microwave driving fields in the first step, and of the optomechanical subsystem and the weak optical pulse in the second step.
The OMM system consists of a magnon mode (e.g., the Kittel mode Kittel ) and an optical cavity mode, and both couple to a mechanical mode via the magnetostrictive and radiation-pressure interactions, respectively, as depicted in Fig. 1(a). The magnomechanical system can be a YIG micro bridge structure PRAp , and the optical cavity can be formed by attaching a small highly-reflective mirror pad to the surface of the micro bridge LPR23 ; SG . Another promising system could be the opto- and magnomechanical systems coupled via direct physical contact Dong22 . The YIG crystal is placed inside a uniform bias magnetic field and further driven by a microwave field applied via, e.g., a loop antenna. The magnon mode is then activated and coupled to the mechanical displacement via the magnetostrictive interaction, and the latter further couples to the optical cavity via radiation pressure. We consider the situation where the mechanical resonance frequency is much lower than the magnon frequency, such that the magnomechanical coupling is a dispersive type Zuo . The Hamiltonian of the three-mode OMM system reads
$H/\hbar=\sum_{j=m,b,c}\omega_{j}j^{\dagger}j+\left(G_{0}m^{\dagger}m-g_{0}c^{\dagger}c\right)\left(b+b^{\dagger}\right)+H_{\rm dri},$ (1)
where $j$ ($j^{\dagger}$), $j=m,b,c$, are the annihilation (creation) operators of the magnon, mechanical, optical cavity modes, respectively, and $\omega_{j}$ are their corresponding frequencies. $G_{0}$ ($g_{0}$) is the bare magnomechanical (optomechanical) coupling strength, and $H_{\rm dri}$ represents the driving Hamiltonian, which is different and will be specified in the following two steps.
In the first step, we aim to squeeze the mechanical motion by driving the magnon with two microwave fields (Fig. 1(b)). Due to the absence of a strong laser driving field to enhance the optomechanical coupling, the effective optomechanical coupling is very weak. The optical cavity is essentially decoupled from the driven magnomechanical system. Therefore, the tripartite OMM system reduces to an effective two-mode system with the Hamiltonian given by
$\begin{split}H_{\rm MM}/\hbar&=\omega_{m}m^{\dagger}m+\omega_{b}b^{\dagger}b+G_{0}m^{\dagger}m\left(b+b^{\dagger}\right)\\ &+\left[\left(\Omega_{+}e^{-i\omega_{+}t}+\Omega_{-}e^{-i\omega_{-}t}\right)m^{\dagger}-\mathrm{H}.\mathrm{c}.\right],\end{split}$ (2)
which indicates that the magnon mode is driven by two microwave fields at the frequencies $\omega_{\pm}=\omega_{m}\pm\omega_{b}$, and $\Omega_{\pm}$ denote the associated magnon-drive coupling strength, i.e., the Rabi frequency JL18 .
Incorporating the dissipation and input noise terms, we obtain the following quantum Langevin equations (QLEs) of the magnomechanical system:
$\begin{split}\dot{m}&=\left(-i\omega_{m}-\frac{\kappa_{m}}{2}\right)m-iG_{0}m\left(b+b^{\dagger}\right)\\ &-i\left(\Omega_{+}e^{-i\omega_{+}t}+\Omega_{-}e^{-i\omega_{-}t}\right)+\sqrt{\kappa_{m}}m_{in},\\ \dot{b}&=\left(-i\omega_{b}-\frac{\kappa_{b}}{2}\right)b-iG_{0}m^{\dagger}m+\sqrt{\kappa_{b}}b_{in},\end{split}$ (3)
where $\kappa_{m}$ and $\kappa_{b}$ are the dissipation rates of the magnon and mechanical modes, respectively, and $O_{in}$ ($O=m,b$) denote the input noises of the two modes, and their nonzero correlation functions are $\langle O_{in}^{\dagger}(t)O_{in}(t^{\prime})\rangle=N_{O}(\omega_{O})\delta(t-t^{\prime})$ and $\langle O_{in}(t)O_{in}^{\dagger}(t^{\prime})\rangle=\left[N_{O}(\omega_{O})+1\right]\delta(t-t^{\prime})$, with $N_{O}(\omega_{O})=\left[{\rm exp}(\hbar\omega_{O}/k_{B}T)-1\right]^{-1}$ being the equilibrium mean thermal magnon/phonon number and $T$ being the bath temperature.
[…]
$\begin{split}m_{\pm}=\frac{\Omega_{\pm}}{\omega_{\pm}-\omega_{m}+i\frac{\kappa_{m}}{2}}.\end{split}$ (4)
We also obtain the linearized QLEs for the quantum fluctuations by neglecting small second-order fluctuation terms, which, in the interaction picture with respect to $\hbar\omega_{m}m^{\dagger}m+\hbar\omega_{b}b^{\dagger}b$, are given by
$\begin{split}\delta\dot{b}&=-\frac{\kappa_{b}}{2}\delta b-i\left(G_{-}+G_{+}e^{2i\omega_{b}t}\right)\delta m\\ &-i\left(G_{+}+G_{-}e^{2i\omega_{b}t}\right)\delta m^{\dagger}+\sqrt{\kappa_{b}}b_{in},\\ \delta\dot{m}&=-\frac{\kappa_{m}}{2}\delta m-i\left(G_{-}+G_{+}e^{-2i\omega_{b}t}\right)\delta b\\ &-i\left(G_{+}+G_{-}e^{2i\omega_{b}t}\right)\delta b^{\dagger}+\sqrt{\kappa_{m}}m_{in}.\end{split}$ (5)
Here, $G_{\pm}=G_{0}m_{\pm}$ are the enhanced magnomechanical coupling strengths due to the two strong driving fields at frequencies $\omega_{\pm}$, which can be set real by adjusting the phases of the drive fields to have real $m_{\pm}$. Due to the time-dependent fast oscillating terms in Eq. (5), the above QLEs are difficult to solve. However, under the condition of $\kappa_{b},\kappa_{m},G_{\pm}\ll\omega_{b}$ JL15 , the rotating-wave approximation can be made by neglecting the fast oscillating terms. Consequently, we obtain
$\begin{split}\delta\dot{b}&=-\frac{\kappa_{b}}{2}\delta b-i\left(G_{-}\delta m+G_{+}\delta m^{\dagger}\right)+\sqrt{\kappa_{b}}b_{in},\\ \delta\dot{m}&=-\frac{\kappa_{m}}{2}\delta m-i\left(G_{-}\delta b+G_{+}\delta b^{\dagger}\right)+\sqrt{\kappa_{m}}m_{in},\end{split}$ (6)
which indicates that the drive field at $\omega_{-}$ is responsible for cooling the mechanical motion by activating the magnomechanical beamsplitter-type interaction (anti-Stokes process), while the drive field at $\omega_{+}$ activates the parametric down-conversion interaction (Stokes process). The proper combination of these two interactions can induce a squeezed mechanical mode, which we show later. To maintain the system stability, the strength of the anti-Stokes scattering should be stronger than that of the Stokes scattering, i.e., $G_{-}>G_{+}$.
Figure 2: Wigner function of the squeezed mechanical mode. The dashed circle corresponds to vacuum fluctuation. The parameters are provided in the main text.
Equation (6) can be rewritten in terms of the quadrature fluctuations, which can be cast in the following matrix form
$\begin{split}\dot{u}(t)=Au(t)+n(t),\end{split}$ (7)
where $u(t)=[\delta X_{m}(t),\,\,\delta Y_{m}(t),\,\,\delta X_{b}(t),\,\,\delta Y_{b}(t)]^{T}$ denotes the vector of the quadrature fluctuations, $n(t)=[\sqrt{\kappa_{m}}X_{m}^{in}(t),\sqrt{\kappa_{m}}Y_{m}^{in}(t),\sqrt{\kappa_{b}}X_{b}^{in}(t),\sqrt{\kappa_{b}}Y_{b}^{in}(t)]^{T}$ is the vector of the input noises, and the quadratures are defined as $X_{O}=\frac{1}{\sqrt{2}}(O+O^{\dagger})$ and $Y_{O}=\frac{i}{\sqrt{2}}(O^{\dagger}-O)$, and $\delta X_{O}$ and $\delta Y_{O}$ are the corresponding fluctuations. Similarly, the associated input noise operators $X_{O}^{in}$ and $Y_{O}^{in}$ can be defined. The drift matrix $A$ is given by
$A=\begin{pmatrix}-\frac{\kappa_{m}}{2}&0&0&G_{+}+G_{-}\\ 0&-\frac{\kappa_{m}}{2}&G_{+}-G_{-}&0\\ 0&G_{+}-G_{-}&-\frac{\kappa_{b}}{2}&0\\ G_{+}+G_{-}&0&0&-\frac{\kappa_{b}}{2}\end{pmatrix}$. (8)
Owing to the linearized dynamics and Gaussian input noises, the steady state of the system is a two-mode Gaussian state, which is fully characterized by a $4\times 4$ covariance matrix (CM) $V$, with its entries defined as $V_{ij}=\langle u_{i}(t)u_{j}(t)+u_{j}(t)u_{i}(t)\rangle/2$ ($i,j=1,2$). The steady-state CM can be obtained by directly solving the Lyapunov equation DV07
$\begin{split}AV+VA^{T}=-D,\end{split}$ (9)
where $D=$ diag[$\big[\kappa_{m}(N_{m}+\frac{1}{2}),\kappa_{m}(N_{m}+\frac{1}{2}),\kappa_{b}(N_{b}+\frac{1}{2}),\kappa_{b}(N_{b}+\frac{1}{2})\big]$ is the diffusion matrix and defined via $D_{ij}\delta(t-t^{\prime})=\langle n_{i}(t)n_{j}(t^{\prime})+n_{j}(t^{\prime})n_{i}(t)\rangle/2$.
For our two-mode Gaussian state, the CM can be expressed in the form
$V\equiv\begin{pmatrix}V_{m}&V_{\rm mb}\\ V_{\rm mb}^{T}&V_{b}\end{pmatrix}$, (10)
and the Wigner function of the mechanical mode can be achieved via RS87
$\begin{split}W(u_{b})=\frac{1}{2\pi\sqrt{{\rm det}V_{b}}}{\rm exp}\left(-\frac{u^{T}_{b}V_{b}^{-1}u_{b}}{2}\right),\end{split}$ (11)
with $u_{b}=(\delta X_{b},\delta Y_{b})^{T}$.
[…]
Figure 3: (a) Degree of squeezing $S$ (dB) of the mechanical mode versus $G_{+}/G_{-}$ for various temperatures. We fix the coupling $G_{-}/2\pi=0.1$ MHz and vary $G_{+}$. (b) Degree of squeezing $S$ (dB) versus bath temperature $T$ and the coupling $G_{-}$. The coupling $G_{+}$ is optimized for each value of $G_{-}$, and the other parameters are as those in Fig. 2.
In Fig. 3(a), we plot the degree of squeezing $S$ (in units of dB) as a function of the ratio $G_{+}/G_{-}$ for various bath temperatures, and $S$ is defined as $S=-10{\rm log}10[V_{\rm min}/V_{\rm vac}]$, with $V_{\rm min}$ being the minimum variance of the mechanical quadrature and $V_{\rm vac}$ being the variance of the vacuum fluctuation. Clearly, there is an optimal ratio $G_{+}/G_{-}$, which decreases as the temperature rises. This is due to the fact that, on the one hand, the difference between the couplings $G_{+}$ and $G_{-}$ should be sufficiently large in order to efficiently cool the mechanical mode; on the other hand, almost equal couplings $G_{+}\simeq G_{-}$ is required to have an extremely large squeezing Jie17 . The trade-off between these two opposite trends results in an optimal ratio $G_{+}/G_{-}$. As the temperature rises, a larger cooling rate is needed to cool the mechanical mode with more thermal noise, which thereby leads to a reduced optimal ratio and a reduced degree of squeezing. In Fig. 3(b), we plot the degree of squeezing $S$ versus the coupling strength $G_{-}$ and the temperature $T$. It tells that the maximum degree of squeezing increases with the coupling $G_{-}$ (with $G_{+}$ being optimized) and decreases with the temperature, in consistency with the results of Fig. 3(a) for a fixed $G_{-}$.
III PHONON SUBTRACTION via optical pulses
Once the mechanical mode is prepared in a squeezed state (ideally a squeezed vacuum state, but practically a squeezed thermal state due to the added thermal noise), we switch off the two microwave drive fields and then send an optical pulse to the optomechanical subsystem to implement the phonon subtraction operation, which we show below can yield a mechanical cat-like state.
Since the magnon driving fields are turned off, the effective magnomechanical coupling becomes very weak. The system, under an optical driving field, then becomes an effective two-mode system with an optomechanical Hamiltonian, given by
$\begin{split}H_{\rm OM}/\hbar=\omega_{c}c^{\dagger}c+\omega_{b}b^{\dagger}b-g_{0}c^{\dagger}c\left(b^{\dagger}+b\right)-E\left(e^{-i\omega_{0}t}a^{\dagger}-\mathrm{H.c.}\right),\end{split}$ (12)
where $E=\sqrt{\kappa_{c}P_{0}/(\hbar\omega_{0})}$ denotes the coupling strength between the optical cavity and the driving field, with $P_{0}$ ($\omega_{0}$) being the power (frequency) of the driving field and $\kappa_{c}$ being the cavity decay rate.
Similarly, we linearize the dynamics following the procedures in Sec. II. When the drive field is red-detuned from the cavity by $\Delta\equiv\omega_{c}-\omega_{0}=\omega_{b}$, and in the resolved sideband limit $\omega_{b}\gg\kappa_{c}$, we derive the following linearized QLEs for the quantum fluctuations $\{\delta c$, $\delta b\}$, which, in the interaction picture with respect to $\hbar\omega_{0}c^{\dagger}c$, are given by JL21
$\begin{split}\delta\dot{c}&=(-i\Delta-\frac{\kappa_{c}}{2})\delta c+iG_{c}\delta b+\sqrt{\kappa_{c}}c_{in},\\ \delta\dot{b}&=(-i\omega_{b}-\frac{\kappa_{b}}{2})\delta b+iG_{c}\delta c+\sqrt{\kappa_{b}}b_{in},\end{split}$ (13)
where $c_{in}$ is the input noise of the cavity, and the effective optomechanical coupling $G_{c}=g_{0}\langle c\rangle$, with $\langle c\rangle=\frac{iE}{i\Delta+\frac{\kappa_{c}}{2}}\simeq\frac{E}{\omega_{b}}$. For simplicity, we consider a flattop pulse such that $G_{c}$ is constant during the pulse. Note that in getting Eq. (13), we have neglected the frequency shift induced by the optomechanical interaction as it is typically much smaller than $\Delta=\omega_{b}$.
We consider the pulse duration to be much shorter than the mechanical lifetime, such that the mechanical dissipation is negligible during the pulse. The corresponding QLEs, in the interaction picture with respect to $\hbar\omega_{c}c^{\dagger}c+\hbar\omega_{b}b^{\dagger}b$, are given by
$\begin{split}\delta\dot{c}&=-\frac{\kappa_{c}}{2}\delta c+iG_{c}\delta b+\sqrt{\kappa_{c}}c_{in},\\ \delta\dot{b}&=iG_{c}\delta c,\end{split}$ (14)
We use weak pulses, yielding $G_{c}\ll\kappa_{c}$, which allows us to adiabatically eliminate the cavity, and obtain $\delta c=i\frac{2G_{c}}{\kappa_{c}}\delta b+\frac{2}{\sqrt{\kappa_{c}}}c_{in}$. By using the input-output relation $c_{out}=\sqrt{\kappa_{c}}\delta c-c_{in}$, we get
$\begin{split}c_{out}&=i\sqrt{2G}\delta b+c_{in},\\ \delta\dot{b}&=-G\delta b+i\sqrt{2G}c_{in},\end{split}$ (15)
where $G=2G_{c}^{2}/\kappa_{c}$. By further introducing the temporal modes associated with the pulse Hammerer
$\begin{split}C_{in}(t)&=\sqrt{\frac{2G}{e^{2Gt}-1}}\int_{0}^{t}dt^{\prime}e^{Gt^{\prime}}c_{in}(t^{\prime}),\\ C_{out}(t)&=\sqrt{\frac{2G}{1-e^{-2Gt}}}\int_{0}^{t}dt^{\prime}e^{-Gt^{\prime}}c_{out}(t^{\prime}),\end{split}$ (16)
which satisfy the commutation relation $C_{k},C_{k}^{\dagger}]=1$ ($k=in,out$), we achieve
$\begin{split}C_{out}(t)&=e^{-Gt}C_{in}(t)+i\sqrt{1-e^{-2Gt}}b(0),\\ \delta b(t)&=e^{-Gt}\delta b(0)+i\sqrt{1-e^{-2Gt}}C_{in}(t).\end{split}$ (17)
From Eq. (17), we can extract a propagator $U(t)$, satisfying $C_{out}(t)=U(t)^{\dagger}C_{in}(t)U(t)$ and $\delta b(t)=U(t)^{\dagger}\delta b(0)U(t)$, given by JL21
$\begin{split}U=e^{i\tan\theta C_{in}^{\dagger}b}\cos\theta^{-(C_{in}^{\dagger}C_{in}-b^{\dagger}b)}e^{-i\tan\theta C_{in}b^{\dagger}}\end{split}$ (18)
where $\cos\theta\equiv e^{-Gt}$ and $\tan\theta\equiv\sqrt{e^{2Gt}-1}$.
We first analyze the ideal case where the mechanical mode is prepared in a squeezed vacuum state $|\xi\rangle_{b}=S(\xi)|0\rangle_{b}$, with $S(\xi)=e^{\frac{1}{2}(\xi b^{\dagger 2}-\xi^{*}b^{2})}$ being the squeezing operator and $\xi=re^{i\phi}$, with $r$ being the squeezing parameter and $\phi$ the phase angle, i.e.,
$\begin{split}|\xi\rangle_{b}&=\frac{1}{\sqrt{\cosh r}}\sum_{n=0}^{\infty}(e^{i\phi}\tanh r)^{n}\frac{\sqrt{(2n)!}}{2^{n}n!}|2n\rangle_{b}\\ &=\sqrt[4]{1-\tanh^{2}r}\sum_{n=0}^{\infty}C_{n}(\tanh r)^{n}|2n\rangle_{b},\end{split}$ (19)
with $C_{n}=\left(\frac{e^{i\phi}}{2}\right)^{n}\frac{\sqrt{(2n)!}}{n!}$. The function of the pulse is equivalent to applying the propagator $U(t)$ onto the initial state $|0\rangle_{c}|\xi\rangle_{b}$, where $|0\rangle_{c}$ denotes the cavity is in vacuum. At the end of the pulse, the system evolves to be
$\begin{split}|\psi\rangle_{c,b}&=U|0\rangle_{c}|\xi\rangle_{b}=\sum_{n=0}^{\infty}\frac{(i\tan\theta)^{\textit{n}}}{n!}\left(C_{in}^{\dagger}b\right)^{n}|0\rangle_{c}|\xi^{\prime}\rangle_{b}\\ &\approx|0\rangle_{c}|\xi^{\prime}\rangle_{b}+i\tan\theta|1\rangle_{c}\big(\textit{b}|\xi^{\prime}\rangle_{\textit{b}}\big)-\frac{\tan^{2}\theta}{\sqrt{2}}|2\rangle_{\textit{c}}\left(\textit{b}^{2}|\xi^{\prime}\rangle_{\textit{b}}\right),\end{split}$ (20)
where
$\begin{split}|\xi^{\prime}\rangle_{b}=\sqrt[4]{1-\left(\tanh r\,\cos^{2}\theta\right)^{2}}\sum_{n=0}^{\infty}C_{n}\left(\tanh r\,\cos^{2}\theta\right)^{n}|2n\rangle_{b}.\end{split}$ (21)
In getting the last line of Eq. (20), we have omitted higher-order terms with $n>2$, which is a good approximation when $\tan^{2}\theta\ll 1$, i.e., $Gt\ll 1$. Equation (20) indicates that the subtraction of a single phonon $b|\xi^{\prime}\rangle_{b}$ can be realized when a single anti-Stokes photon in the cavity output field is detected, and the successful probability is approximately $\tan^{2}\theta$. Similarly, the subtraction of two phonons $b^{2}|\xi^{\prime}\rangle_{b}$ can also be achieved, with a much smaller probability $\sim\tan^{4}\theta$, conditioned on the detection of a two-photon state. We note that due to the pulse interaction, the squeezing $r$ of the squeezed vacuum state is slightly reduced, because $\tanh r\rightarrow\tanh r\,\cos^{2}\theta$ as seen in Eq. (21).
In a practical situation, as demonstrated in Sec. II, we prepare a squeezed thermal state due to the nonnegligible thermal noise at a finite temperature, which is reflected in the CM $V_{b}$ of the mechanical mode. To be compatible with the propagator $U(t)$, we transform the CM into a density matrix $\rho_{in,b}$ (Appendix), which acts as the initial state before the optical pulse is applied. Combining the vacuum state of the cavity, the initial state of the optomechanical system is then $\rho_{in}=\rho_{in,b}\otimes|0\rangle\langle 0|_{c}$, which, at the end of the pulse, evolves into the state $\rho_{b,c}=U\rho_{in}U^{\dagger}$, given by MD97 ; MB94
$\begin{split}\rho_{b,c}&=\sum_{n=0}^{\infty}\sum_{m=0}^{\infty}\frac{e^{-i(m-n)\pi/2}}{\sqrt{n!m!}}(-1)^{m+n}|\tan\theta|^{m+n}\\ &\times b^{m}|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}\textit{b}^{\dagger n}\otimes|m\rangle\langle n|_{\textit{c}}.\end{split}$ (22)
We consider the situation where a $n$-photon state is detected in the output field, i.e., taking $m=n$ in Eq. (22), which gives rise to the following state
$\begin{split}\rho_{b,c}&=\sum^{\infty}_{n=0}\frac{(\tan\theta)^{2n}}{n!}b^{n}|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger n}\otimes|\textit{n}\rangle\langle\textit{n}|_{\textit{c}}\\ &\approx|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}\otimes|0\rangle\langle 0|_{\textit{c}}\\ &+\tan^{2}\theta\,b|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger}\otimes|1\rangle\langle 1|_{\textit{c}}\\ &+\frac{\tan^{4}\theta}{2}b^{2}|\cos\theta|^{b^{\dagger}b}\rho_{in,b}|\cos\theta|^{b^{\dagger}b}b^{\dagger 2}\otimes|2\rangle\langle 2|_{\textit{c}},\end{split}$ (23)
where higher-order terms with $n>2$ are neglected because $\tan^{2}\theta\ll 1$. The above equation indicates that a single phonon (two phonons) can be subtracted from the initial squeezed thermal state, i.e., $b\rho_{in,b}b^{\dagger}$ ($b^{2}\rho_{in,b}b^{\dagger 2}$), conditioned on the detection of a single-photon (two-photon) state in the cavity output. Note that the slightly reduced $r$ as revealed in Eq. (21) is not visible because a general mixed state $\rho_{in,b}$ is adopted in Eq. (23).
Figure 4: Wigner function of (a) the single-phonon [(b) the two-phonon] subtracted squeezed mechanical state. The parameters are provided in the text.
In Fig. 4, we plot the Wigner function of the single/two-phonon-subtracted squeezed mechanical state. The phonon subtraction leads to the negativity of the Wigner function, demonstrating the nonclassical nature of the state. Moreover, the Wigner function exhibits a coherent superposition feature with the interference fringes shown around the origin of the phase space, which resembles that of the Schrödinger’s cat state. For this reason, it is called mechanical cat-like state and its fidelity with the cat state increases when more phonons are subtracted from the squeezed state IS20 ; Tan20 , which corresponds to a smaller successful probability and thus longer measurement time. We take the following parameters: optical wavelength $\lambda_{0}=1550$ nm, $\kappa_{c}/2\pi=3$ MHz, $g_{0}/2\pi=2$ kHz, pulse duration $t=30$ ns and power $P_{0}=50$ pW, which yield $G/2\pi=1.8$ kHz and $\tan\theta\approx 0.11$. The probability of subtracting a single phonon (two phonons) is about $\sim 1\%$ ($0.01\%$).
IV Conclusions
We present a two-step scheme to prepare a cat-like state of mechanical motion in an OMM system. By applying a two-tone microwave drive field to the magnomechanical system, the mechanical mode is prepared in a squeezed thermal state in the first step. After turning off the microwave dirves, by further sending a weak red-detuned optical pulse to the optomechanical cavity, $k$ phonons can be subtracted from the squeezed mechanical mode conditioned on the detection of $k$ anti-Stokes photons from the cavity output field, which results in a mechanical cat-like state. More phonons subtracted from the squeezed state leads to a higher fidelity with the cat state, but a smaller successful probability. The work may find applications in macroscopic quantum studies and quantum sensing and in the test of collapse models.
ACKNOWLEDGMENTS
This work was supported by National Key Research and Development Program of China (Grants No. 2024YFA1408900 and No. 2022YFA1405200), Zhejiang Provincial Natural Science Foundation of China (Grant No. LR25A050001), and National Natural Science Foundation of China (Grants No. 12474365 and No. 92265202).
APPENDIX: TRANSFORMATION FROM COVARIANCE MATRIX TO DENSITY MATRIX
In Sec. II, we prepare a squeezed thermal mechanical state characterized by a $2\times 2$ CM $V_{b}$, which acts as the initial state $\rho_{in,b}$ for the subsequent operation of phonon subtraction. To exploit the derived propagator $U(t)$ associated with the pulse, we need to transform the form of the mechanical state from the CM to the density matrix. The squeezed thermal state can be expressed as $\rho_{in,b}=S(\xi)\rho_{th}(\bar{n})S^{\dagger}(\xi)$, where $S(\xi)$ is the squeezing operator, $\xi=re^{i\phi}$, and $\rho_{th}=\Sigma_{n}P_{n}|n\rangle\langle n|$ is the thermal state, $P_{n}=\bar{n}^{n}/(1+\bar{n})^{n+1}$, with $\bar{n}$ being the mean thermal phonon number. The characteristic parameters $r$, $\phi$, and $\bar{n}$ can be extracted from the CM $V_{b}$ via VG06
$\begin{split}r&=\frac{1}{2}{\rm arcosh}\,\frac{{\rm Tr}\,V_{b}}{2\sqrt{{\rm det}\,V_{b}}},\\ \bar{n}&=\frac{\sqrt{{\rm det}\,V_{b}}-1}{2},\\ \tan\phi&=\frac{2V_{12}}{V_{11}-V_{22}},\end{split}$ (A1)
with $V_{j,k}$ ($j,k=1,2$) being the elements of $V_{b}$. For the mechanical state with $V_{b}=\{\{0.045,0.14\},\{0.14,6.28\}\}$ prepared in Sec. II, we get $r=1.25$, $\phi=-0.04$, and $\bar{n}=0.013$.